# The Inference of Fokker-Planck Equations via Transport Maps

### v2

@@ 2.4 Model Framework
Figure 3: Overview of the model architecture. The pullback map $S$ and the the Fokker-Planck equation quantities $\Psi,D$ are parametrized by $\bm{\Theta}_{1}$ and $\bm{\Theta}_{2}$, respectively, and optimized based on data observed at discrete time points.
@@
The KL divergence in (29) decomposes as:
$D_{KL}(\rho\|S^{\#}\eta)=-\sum_{X}\left(\rho\log S^{\#}\eta-\rho\log\rho\right).$ (30)
@@ 3.1.1 Isotropic diffusion with quadratic potential functions
$\Psi(x)=\frac{1}{2}x^{T}\begin{pmatrix}\theta_{\Psi}^{1}&0\\ 0&\theta_{\Psi}^{2}\end{pmatrix}x+\begin{pmatrix}\theta_{\Psi}^{3}\\ \theta_{\Psi}^{4}\end{pmatrix}^{T}.$ (44)
@@
3.1.2 Anisotropic diffusion with double-well potential functions
We next consider dynamics which are governed by anisotropic diffusion and the double-well potential function:
@@ 3.1.2 Anisotropic diffusion with double-well potential functions
$\tilde{D}=\begin{pmatrix}\theta_{D}^{1}&\theta_{D}^{2}\\ \theta_{D}^{3}&\theta_{D}^{4}\end{pmatrix}$. (46)
Since the diffusion tensor must be symmetric, we represent it as $D=\frac{1}{2}(\tilde{D}+\tilde{D}^{T})$. The potential function is then modeled in two ways as in the previous problem, first, with the pre-defined structure:
@@
and second, using a general neural network. The recovered potential functions along with the ground truth are shown in Figure 8 and Figure 9. The structured model clearly recovers the underlying potential function more accurately, while the neural network model, despite being unconstrained, successfully captures the shape of the wells. Table 2 presents the inferred diffusion and potential parameters under the structure-form assumption. The initial and final data samples superposed on the double-well potential function are displayed in Figure 10.
@@
Figure 8: Contour plots of the potential function. Each column shows: (left) the ground truth potential, (middle) reconstructed potential using the structured model, and (right) the reconstructed potential using the neural network model for the two-dimensional anisotropic, double-well potential problem.
@@
Figure 9: Slices of the potential function along each axis: (left) at $x_{2}=0$ and (right) at $x_{1}=0$ for the two-dimensional anisotropic, double-well potential problem.
@@
Figure 10: The initial (left) and the final (right) data samples (blue dots) superposed on the contour plot of the Fokker-Planck potential function for the two-dimensional anisotropic, double-well potential problem.
$\theta_{D}^{1}$ $\theta_{D}^{2}$ $\theta_{D}^{3}$ $\theta_{D}^{4}$ $\theta_{\Psi}^{1}$ $\theta_{\Psi}^{2}$ $\theta_{\Psi}^{3}$ $\theta_{\Psi}^{4}$
Ground truth 0.1 0.0 0.0 0.2 1.0 1.0 -1.5 -1.5
Recovered 0.11 -0.01 -0.01 0.17 1.01 0.91 -1.45 -1.21
@@
Table 2: Ground truth and recovered system parameters for the two-dimensional anisotropic, double-well potential problem.
@@
The neural network representation of the pullback map is obtained by training the model, which enables the computation of the densities by (3). The evolution of the data distribution and the reconstructed probability density functions using two distinct models are illustrated in Figure 11. The probabilistic model in (38) further enables uncertainty quantification. In particular, Bayesian neural networks provide estimates of the mean and variance of the reconstructed density, as shown in Figure 12.
@@
Figure 11: Contour plots of the density at different time instances. The rows correspond to a different time stamp: $t=0.1,0.4,0.7$ (from top to bottom). Each column shows: (left) ground truth density, (middle) reconstructed density using the structured potential model, and (right) reconstructed density using the neural network potential model for the two-dimensional anisotropic, double-well potential problem.
@@
Figure 12: Mean and standard deviation of the reconstructed density at $t=1.0$ generated by bayesian neural networks with a probabilistic architecture. (Left) ground truth distribution, (middle) mean of the reconstructed density, and (right) standard deviation of the reconstructed density obtained using the neural network potential model for the two-dimensional anisotropic, double-well potential problem.
@@ 3.2 A high-dimensional Fokker-Planck problem
In this section, we apply our method to Fokker-Planck dynamics in five-dimensions, demonstrate the convergence of the inferred diffusion tensor components and parameters of the structured potential representation, as well as present projections of the inferred neural network potential surface. In the context of high-dimensional system inference, Chen et al. previously proposed an algorithm using the general Physics-Informed Neural Network framework, which captures both the governing equation and the underlying density function [47]. Their approach involves defining a loss function based on the PDE residual evaluated at collocation points. In similar numerical experiments, they used $100,000$ samples for each of seven time snapshots, imposed the PDE constraint using mini-batches of size $50,000$, and evaluated the residual on $10,000$ points. We show that our method achieves comparable convergence with significantly fewer data points.
@@
$\tilde{D}=\begin{pmatrix}\theta_{D}^{1,1}&\theta_{D}^{1,2}&\theta_{D}^{1,3}&\theta_{D}^{1,4}&\theta_{D}^{1,5}\\ \theta_{D}^{2,1}&\theta_{D}^{2,2}&\theta_{D}^{2,3}&\theta_{D}^{2,4}&\theta_{D}^{2,5}\\ \theta_{D}^{3,1}&\theta_{D}^{3,2}&\theta_{D}^{3,3}&\theta_{D}^{3,4}&\theta_{D}^{3,5}\\ \theta_{D}^{4,1}&\theta_{D}^{4,2}&\theta_{D}^{4,3}&\theta_{D}^{4,4}&\theta_{D}^{4,5}\\ \theta_{D}^{5,1}&\theta_{D}^{5,2}&\theta_{D}^{5,3}&\theta_{D}^{5,4}&\theta_{D}^{5,5}\end{pmatrix},\quad\Psi=-\log\left(e^{-\sum_{i=1}^{5}(x_{i}+\theta_{\Psi}^{1,i})^{2}}+e^{-\sum_{i=1}^{5}(x_{i}+\theta_{\Psi}^{2,i})^{2}}\right).$ (49)
Additionally, we used the neural network representation of the potential. Since the diffusion tensor must be symmetric, we enforce this by expressing it as $D=\frac{1}{2}(\tilde{D}+\tilde{D}^{T})$. We track the convergence of the system parameters (diffusion tensor components and structured potential representation) over iterations, with the results shown in Figure 13 for the structure-informed potential function representation to allow comparison with the ground-truth synthetic data. Comparable results were obtained for the diffusion tensor when the full neural network representation was used for the potential with $12,000$ particles. Importantly, the transport map representation is always with neural networks. The results obtained using the neural network potential representation are illustrated in Figure 14 for convergence of the diffusion tensor, and in Figures 15, 16, 17 , which show two-dimensional slices through the surface that is the relative error in the neural network potential. The ground-truth data on particle positions projected onto the corresponding plane at $t=0.5$ is super-posed on the relative error in the neural network potential surface. The particle color represents the projection distance. While each two-dimensional slice of the ground-truth potential is identical by symmetry, this constraint was not imposed on the structured and neural network representations. The learnt neural network representations differ between two-dimensional slices, because the diffusional motion of $12,000$ particles over $\mathbb{R}^{5}$ provides sparse information and unsymmetric particle distributions to resolve the surface. This results in different relative errors in the neural network potential surface for each slice. However, the minima within the slices are in good approximation. With lower diffusion, the samples would be more localized to the minima.
@@ 4 Discussion and conclusion
In this paper, we have presented a framework for inferring the Fokker-Planck equation underlying the dynamics of an observed data distribution, while simultaneously learning an explicit transport map-based representation of the corresponding probability densities. This method is built upon the Knothe-Rosenblatt rearrangement, a triangular transport map that is both computationally tractable and theoretically well-established.
A key observation is that our approach requires far fewer data points to achieve robust performance, even compared to nontraditional grid-free methods. Empirically, in the absence of prior structural knowledge, approximately $1,000$ samples per time step were needed in two-dimensional settings to achieve a suitable accuracy with respect to ground truth data. Comparable results were obtained using around $3,000$, $7,000$, and $12,000$ samples per time step in three, four, and five dimensions, respectively. When prior structural information is incorporated, the required sample size decreased further; for instance only 3000 samples were needed in five dimensions. Even when integrals are estimated via random sampling, the transport-based representation adaptively assigns greater weight to regions where the data is concentrated, without requiring explicit knowledge of the boundaries of the data distribution. The above reported sample numbers are lower by orders of magnitude when compared with other methods in the literature. We treat the uncertainty introduced by finite data samples in a standard approach via Bayesian neural networks trained using variational inference.

### v3

@@ 2.4 Model Framework
Figure 3: Overview of the model architecture. The pullback map $S$ and the Fokker-Planck equation quantities $\Psi,D$ are parametrized by $\bm{\Theta}_{1}$ and $\bm{\Theta}_{2}$, respectively, and optimized based on data observed at discrete time points.
@@
Given a sample space $X$, the KL divergence in (29) decomposes as:
$D_{KL}(\rho\|S^{\#}\eta)=-\sum_{x\in X}\left(\rho(x)\log(S^{\#}\eta)(x)-\rho(x)\log\rho(x)\right).$ (30)
@@ 3.1.1 Isotropic diffusion with quadratic potential functions
$\Psi(x)=\frac{1}{2}x^{T}\begin{pmatrix}\theta_{\Psi}^{1}&0\\ 0&\theta_{\Psi}^{2}\end{pmatrix}x+\begin{pmatrix}\theta_{\Psi}^{3}\\ \theta_{\Psi}^{4}\end{pmatrix}^{T}x.$ (44)
@@
3.1.2 Anisotropic diffusion with annular potential functions
We next consider dynamics which are governed by anisotropic diffusion and the annular potential function:
@@ 3.1.2 Anisotropic diffusion with annular potential functions
$D=\begin{pmatrix}\theta_{D}^{1}&\theta_{D}^{2}\\ \theta_{D}^{2}&\theta_{D}^{3}\end{pmatrix}$ (46)
for symmetric diffusion tensors. The potential function is then modeled in two ways as in the previous problem, first, with the pre-defined structure:
@@
and second, using a general neural network. The recovered potential functions along with the ground truth are shown in Figure 8 and Figure 9. The structured model clearly recovers the underlying potential function more accurately, while the neural network model, despite being unconstrained, successfully captures the shape of the wells. Figure 10 shows the mean and standard deviation of the Bayesian neural network-learned potential plotted in comparison with the ground truth. Table 2 presents the inferred diffusion and potential parameters under the structure-form assumption. The initial and final data samples superposed on the annular potential function are displayed in Figure 11.
@@
Figure 8: Contour plots of the potential function. Each column shows: (left) the ground truth potential, (middle) reconstructed potential using the structured model, and (right) the reconstructed potential using the neural network model for the two-dimensional anisotropic, annular potential problem.
@@
Figure 9: Slices of the potential function along each axis: (left) at $x_{2}=0$ and (right) at $x_{1}=0$ for the two-dimensional anisotropic, annular potential problem.
Figure 10: Slices of the potential function reconstructed by the Bayesian framework. The profiles are shown along each axis: (left) at $x_{2}=0$ and (right) at $x_{1}=0$ for the two-dimensional anisotropic, annular potential problem. The dashed lines represent the mean, while the shaded regions illustrate the standard deviation.
@@
Figure 11: The initial (left) and the final (right) data samples (blue dots) superposed on the contour plot of the Fokker-Planck potential function for the two-dimensional anisotropic, annular potential problem.
$\theta_{D}^{1}$ $\theta_{D}^{2}$ $\theta_{D}^{3}$ $\theta_{\Psi}^{1}$ $\theta_{\Psi}^{2}$ $\theta_{\Psi}^{3}$ $\theta_{\Psi}^{4}$
Ground truth 0.1 0.0 0.2 1.0 1.0 -1.5 -1.5
Recovered 0.11 -0.01 0.17 1.01 0.91 -1.45 -1.21
@@
Table 2: Ground truth and recovered system parameters for the two-dimensional anisotropic, annular potential problem.
@@
The neural network representation of the pullback map is obtained by training the model, which enables the computation of the densities by (3). The evolution of the data distribution and the reconstructed probability density functions using two distinct models are illustrated in Figure 12. The probabilistic model in (38) further enables uncertainty quantification. In particular, Bayesian neural networks provide estimates of the mean and variance of the reconstructed density, as shown in Figure 13. As might be expected, the regions of high variance in this plot correspond to higher values of the residual component of the objective function (22), defined as $\left|\rho v+\rho\nabla\widetilde{\Psi}+\widetilde{D}\nabla\rho\right|^{2}$, as shown in Figure (14).
@@
Figure 12: Contour plots of the density at different time instances. The rows correspond to a different time stamp: $t=0.1,0.4,0.7$ (from top to bottom). Each column shows: (left) ground truth density, (middle) reconstructed density using the structured potential model, and (right) reconstructed density using the neural network potential model for the two-dimensional anisotropic, annular potential problem.
@@
Figure 13: Mean and standard deviation of the reconstructed density at $t=1.0$ generated by bayesian neural networks with a probabilistic architecture. (Left) ground truth distribution, (middle) mean of the reconstructed density, and (right) standard deviation of the reconstructed density obtained using the neural network potential model for the two-dimensional anisotropic, annular potential problem.
Figure 14: The squared residual norm evaluated after Bayesian training for the two-dimensional anisotropic, annular potential problem. The mapping illustrates regions where the constraints are less strictly satisfied, providing a visual proxy for the total predictive uncertainty across the domain.
Remark: This work utilizes a Bayesian framework, specifically Bayesian Neural Networks (BNNs), to provide robust quantification of the uncertainty arising due to sparseness of the data. It is important to clarify that the observed residuals integrate uncertainty from three interconnected components: the transport map, the potential function, and the diffusion tensor. Due to the high degree of coupling between these variables, a rigorous causal attribution of uncertainty of each individual source remains a non-trivial challenge. We also note that variational inference, which has been employed in this work, allows a formal decomposition into the aleatoric and epistemic components of uncertainty using approaches such as Variational Uncertainty Decomposition [47]. Such a deeper treatment of uncertainty is reserved as an important objective for future study to further enhance model interpretability.
@@ 3.2 A high-dimensional Fokker-Planck problem
In this section, we apply our method to Fokker-Planck dynamics in five-dimensions. A primary motivating example is the analysis of experimental data on cell dynamics of molecular signaling and migration. Previous studies have demonstrated that cell migration can be modeled using variants of the Fokker-Planck equation [48]. Cell signaling involves distinct chemical reporters interacting with each other and the spatial position variables of migration. The number of these reporters can range from one to tens. With such an extension in mind, we demonstrate the convergence of the inferred diffusion tensor components and parameters of the structured potential representation, as well as present projections of the inferred neural network potential surface. In the context of high-dimensional system inference, Chen et al. previously proposed an algorithm using the general Physics-Informed Neural Network framework, which captures both the governing equation and the underlying density function [49], also in five dimensions. Their approach involves defining a loss function based on the PDE residual evaluated at collocation points. In similar numerical experiments, they used $100,000$ samples for each of seven time snapshots, imposed the PDE constraint using mini-batches of size $50,000$, and evaluated the residual on $10,000$ points. We show that our method achieves comparable convergence with significantly fewer data points.
@@
$D=\begin{pmatrix}\theta_{D}^{1,1}&\theta_{D}^{1,2}&\theta_{D}^{1,3}&\theta_{D}^{1,4}&\theta_{D}^{1,5}\\ \theta_{D}^{1,2}&\theta_{D}^{2,2}&\theta_{D}^{2,3}&\theta_{D}^{2,4}&\theta_{D}^{2,5}\\ \theta_{D}^{1,3}&\theta_{D}^{2,3}&\theta_{D}^{3,3}&\theta_{D}^{3,4}&\theta_{D}^{3,5}\\ \theta_{D}^{1,4}&\theta_{D}^{2,4}&\theta_{D}^{3,4}&\theta_{D}^{4,4}&\theta_{D}^{4,5}\\ \theta_{D}^{1,5}&\theta_{D}^{2,5}&\theta_{D}^{3,5}&\theta_{D}^{4,5}&\theta_{D}^{5,5}\end{pmatrix},\quad\Psi=-\log\left(e^{-\sum_{i=1}^{5}(x_{i}+\theta_{\Psi}^{1,i})^{2}}+e^{-\sum_{i=1}^{5}(x_{i}+\theta_{\Psi}^{2,i})^{2}}\right).$ (49)
Additionally, we used the neural network representation of the potential. Since the diffusion tensor must be symmetric, we enforce this by expressing it as $D=\frac{1}{2}(\tilde{D}+\tilde{D}^{T})$. (Alternatively, one may parametrize only the lower triangular portion of the matrix to improve computational efficiency.) We track the convergence of the system parameters (diffusion tensor components and structured potential representation) over iterations, with the results shown in Figure 15 for the structure-informed potential function representation to allow comparison with the ground-truth synthetic data. Comparable results were obtained for the diffusion tensor when the full neural network representation was used for the potential with $12,000$ particles. Importantly, the transport map representation is always with neural networks. The results obtained using the neural network potential representation are illustrated in Figure 16 for convergence of the diffusion tensor, and in Figures 17, 18, 19 , which show two-dimensional slices through the surface that is the relative error in the neural network potential. The ground-truth data on particle positions projected onto the corresponding plane at $t=0.5$ is super-posed on the relative error in the neural network potential surface. The particle color represents the projection distance. While each two-dimensional slice of the ground-truth potential is identical by symmetry, this constraint was not imposed on the structured and neural network representations. The learnt neural network representations differ between two-dimensional slices, because the diffusional motion of $12,000$ particles over $\mathbb{R}^{5}$ provides sparse information and unsymmetric particle distributions to resolve the surface. This results in different relative errors in the neural network potential surface for each slice. However, the minima within the slices are in good approximation. With lower diffusion, the samples would be more localized to the minima.
@@ 4 Discussion and conclusion
In this paper, we have presented a framework for inferring the Fokker-Planck equation underlying the dynamics of an observed data distribution, while simultaneously learning an explicit transport map-based representation of the corresponding probability densities. This method is built upon the Knothe-Rosenblatt rearrangement, a triangular transport map that is both computationally tractable and theoretically well-established. In this regard, we note that the variable ordering in the Rosenblatt transformation can influence the results. Generally, there is no universal criterion for determining an optimal coordinate sequence. The ideal permutation is often problem-dependent and requires domain-specific insights or additional structural assumptions. While developing systematic strategies for identifying optimal ordering is a compelling direction for future research, it remains beyond the scope of this study.
A key observation is that our approach requires far fewer data points to achieve robust performance, even compared to nontraditional grid-free methods. Empirically, in the absence of prior structural knowledge, approximately $1,000$ samples per time step were needed in two-dimensional settings to achieve a suitable accuracy with respect to ground truth data. Comparable results were obtained using around $3,000$, $7,000$, and $12,000$ samples per time step in three, four, and five dimensions, respectively, with motivation coming from cell dynamics in high dimensions. See Figure 20, where we have included a fit of $\log\,N$ (samples) versus dimension $d$, which demonstrates $N=152.93\times d^{2.71}$ scaling.
Figure 20: Empirical assessment of sample requirements. The plot illustrates the number of training samples required to achieve a stable reconstruction, demonstrating the model’s scaling behavior. The empirical scaling is of the form $N=152.93\times d^{2.71}$.
When prior structural information is incorporated, the required sample size decreased further; for instance only $3,000$ samples were needed in five dimensions. Even when integrals are estimated via random sampling, the transport-based representation adaptively assigns greater weight to regions where the data is concentrated, without requiring explicit knowledge of the boundaries of the data distribution. The above reported sample numbers are lower by orders of magnitude when compared with other methods in the literature. As noted in Section 3.2, the transport map-based approach is more efficient by at least an order of magnitude in comparison with the work of Chen et al. [50].
We treat the uncertainty introduced by finite data samples in a standard approach via Bayesian neural networks trained using variational inference. The framework of variational inference allows a decomposition into the aleatoric and epistemic components of uncertainty using approaches such as Variational Uncertainty Decomposition [47], which we reserve for future work.
@@
Code
Our code is available at github.com/mechanoChem/transportMaps